OpenStudy (anonymous):

Find the slope of the points (-2,8) (-6,5)

OpenStudy (annon):

slope formula (y2-y1)/(x2-x1) = (5-8)/(-6 - -2) = 3/4

OpenStudy (anonymous):

annon is correct: m=3/4

OpenStudy (anonymous):

oki had that one thanx